Daphne and Niles are both driving west on Route 66. At 4:00pm, Daphne is 5 miles west of Niles. A little later, Niles passes Daphne. If Niles is driving at 55 mph and Daphne at 50 mph, at what times does Niles pass Daphne?
Answer by stanbon(75887)

Daphne and Niles are both driving west on Route 66. At 4:00pm, Daphne is 5 miles west of Niles. A little later, Niles passes Daphne. If Niles is driving at 55 mph and Daphne at 50 mph, at what times does Niles pass Daphne?
--------------
Draw the picture letting the position of the two cars at 4:00 be the
starting point.
Daphne is 5 miles ahead of Niles at 4:00
------------------------
Niles DATA:
distance = x+5 miles ; rate = 55 mph ; time = (x+5)/55
--------------
Daphne DATA:
distance = x miles ; rate = 50 mph; time = x/50
--------------
EQUATION:
time = time
(x+5)/55 = x/50
50x+250 = 55x
5x = 250
x = 50 miles
Time = x/50 = 50/50 = 1 hr
They are at the same point at 5:00pm
===============
Cheers,
Stan H.